Real GDP    Consumption Savings    Investment

2000 2000    0 1500

4000 3500    500    1500

6000    5000    1000    1500

8000    6500    1500 1500 (S=I at equilibrium level)

10000    8000    2000    1500

12000 9500 2500 1500

At equilibrium real​ GDP, this​ economy's average propensity to save is
Answer #1
Real GDP Consumption Savings Investment C + I
2000 2000 0 1500 3500
4000 3500 500 1500 5000
6000 5000 1000 1500 6500
8000 6500 1500 1500 8000
10000 8000 2000 1500 9500
12000 9500 2500 1500 11000

At equilibrium: C + I = Real GDP

Equilibrium Real GDP = 8000

Average propensity to save = Savings/Real GDP = 1500/8000 = 0.1875